OpenStudy (miasan5653):

Hey, Can someone help on my little brothers math test. He got it wrong and I don't understand why.

OpenStudy (miasan5653):

In the month of October, the population of a species of fish in a river increased by a factor of 1.05 every day. The function below shows the number of fish in the river f(x) after x days: f(x) = 100(1.05)x Which of the following is a reasonable domain for the function? 0 ≤ x ≤ 100 0 ≤ x ≤ 31 All positive integers greater than 100 All positive integers less than 100

OpenStudy (otherworldly):

which did he pick in this one

OpenStudy (miasan5653):

C

OpenStudy (mathmale):

Mia: If your brother is at this level of math, he is old enough to have his own OpenStudy account. If that is the case, please encourage him to sign up directly and get help directly. Kind of you to want to help, tho'

OpenStudy (nieves):

It should be b, though

OpenStudy (miasan5653):

He doesn't really care. I just tried to help him and am now stuck. Why would it be B

OpenStudy (mathmale):

Kindly do NOT deal in direct answers.

OpenStudy (nieves):

sorry

OpenStudy (miasan5653):

Is the domain dependent on the number of days of the month then?

OpenStudy (mathmale):

f(x) = 100(1.05)x should be written as f(x) = 100(1.05)^x. The " ^ " denotes exponentiation. What does "domain" mean to you? Have to understand that concept first.

OpenStudy (miasan5653):

I get it now. I wasn't thinking it through. I just assumed it meant number of fish based off of the other answers. Thanks

OpenStudy (mathmale):

"Domain" here refers to ACCEPTABLE values of the input variable, x. In other words, find the set of numbers x for which the function is definited.